Grade 11General Physics

The ratio of orbital velocoty to escape velocity is

Hussayn Khan
8 Years agoGrade 11

2 Answers

Amritpal Singh Saini
8 Years ago
Orbital velocity is given by √gR , where R is the radius of earth and escape velocity is given by √2gR Thus the ratio of orbital velocity to escape velocity is 1/√2 (or 0.7070)
Arun
8 Years ago
Orbital speed at the surface is given by 

GM/R^2 = V^2 / R 

V = √ G M / R 

For escape velocity 

1/2 . m . V(e)^2 = GMm / R 

V(e) = √ 2 . G . M / R 

V(e) / V = √2 

( This is the same for any satellite around any planet or star )